Hi need some help Calculating the Liquidity, solvency and profitability of Marriott Intercontinental with the Financial Statement of Year 2012. Please, I would appreciate a brief description of how was calculated everything to understand the exercise.

Liquidity Working capital Current ratio Current cash debt coverage Inventory turnover Days in inventory Accounts receivable turnover Average collection period Current assets-Current liabilities Current assets Current liabilities Net cash provided by operating activities Average current liabilities Cost of goods sold Average inventory 365 days Inventory turnover Net credit sales Average net accounts receivable 365 days receivable turnover Solvency Debt to assets ratio Cash debt coverage Times interest earned Total assets Net cash provided Average ing activities Net income+ Interest expense + Tax expense Interest expense Net cash provided by Capital operating activies xpenditures dividends Cash Free cash flow Profitability Net income Preferred dividends Earnings per share Stock price per share Earnings per share Gross proft Net sales Gross profit rate Net income Net sales Profit margin Net income Average total assets Return on assets Net sales Average total assets Asset turnover Cash dividends declared on common stock Net income Payout ratio Return on common stockholders equity Net income-Preferreddvidends - Average common stockholders equity
media%2Fdfd%2Fdfd6663e-6c11-4e9d-9ce9-e2
media%2F56b%2F56b4243f-752c-4d73-9663-c1
media%2Fc9c%2Fc9c7f692-825e-49bc-88a0-11
media%2Fb2b%2Fb2b79a4d-87cc-4eff-a788-7c
media%2F661%2F66183a48-a406-4da7-b2da-35
media%2Ff2e%2Ff2ec5cd6-c8ab-49d1-a236-ae
Answer #1

Liquidity

Working Capital = 1475-2773 = -1298
Current Ratio= 1475/2773 =0.53
Current cash debt coverge = 989/2773 = 0.36
Inventory Turnover= Not Applicable- Inventory info not provided
Days in Inventory= Not Applicable- Inventory info not provided
Accounts Receivable turnover= 989/1028 =0.96
Average Collection Period = 365/0.96 =381 days

Solvency

Debt to assets Ratio = 898/1989 =0.45
Cash debt coverage = 989/1989 =0.50
Times Interest Earned = 571+137+278/137 =7.19
Free Cash Flow = 989-437-191 =361

Profitability

Earnings Per Share= 571/322 =1.77
Price Earnings Ratio = Market price of share not given in question
Gross Profit rate = 940/11814 =0.08
Net Profit rate = 571/11814 =0.05